OpenStudy (anonymous):

Can you please help me I do not know how to simplify the infinite product of (1+x)(1+x2)(1+x4)(1+x8)(1+x16).., given |x| < 1. Please help....

ganeshie8 (ganeshie8):

consider the partial products : n=1 : 1+x n=2 : (1+x)(1+x^2) = 1+x+x^2 + x^3 ...the pattern continues n=k : 1+x+x^2 + ... x^(2^k-1)

ganeshie8 (ganeshie8):

express it as partial sum for n=k, and take the limit \(k \to \infty \)

OpenStudy (anonymous):

hmmmm i'm not following you

OpenStudy (somy):

maybe some binomial theorm will help ?

OpenStudy (somy):

not binomial itself , the inverse of it

ganeshie8 (ganeshie8):

for n = k : \[\large 1+x+x^2 + \cdots x^{2^k-1} = \dfrac{1-x^{2^k}}{1-x}\]

OpenStudy (anonymous):

ok...

ganeshie8 (ganeshie8):

does above sum looks okay ? :)

OpenStudy (somy):

|x|<1 is givven for a reason :D

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it is correct

ganeshie8 (ganeshie8):

the key thing to notice in the start is this : (a+b)(c+d) gives you 2*2 = 4 = 2^2 terms (a+b)(c+d)(e+f) gives you 2*2*2 = 8 = 2^3 terms

ganeshie8 (ganeshie8):

as you increase the number of binomials in ur product, the terms in expansion increases in powers of 2

OpenStudy (anonymous):

oooookkkkkk.....

ganeshie8 (ganeshie8):

take the limit for the infinite product : \[\large 1+x+x^2 + \cdots x^{2^k-1} + \cdots = \lim \limits_{k\to \infty } \dfrac{1-x^{2^k}}{1-x}\]

OpenStudy (anonymous):

ohhhh i see. Its getting a bit clearer now

ganeshie8 (ganeshie8):

since |x| < 1, \(x^{2^k} \to 0 \) as \(k \to \infty \) @Somy

OpenStudy (somy):

i roger that xD i only need to get back to my acount to loo smart in math :P

OpenStudy (anonymous):

ok i think i can n from here but is it ok if you wait and check if i got the right answer?

ganeshie8 (ganeshie8):

sure :)

ganeshie8 (ganeshie8):

how did u manage to get into all these accounts ? are you the one using Kainui's account also ? :o

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i believe the answer is\[\frac{ 1 }{ 1-x }\]

OpenStudy (anonymous):

am i correct?

ganeshie8 (ganeshie8):

Looks good ^^

OpenStudy (anonymous):

you sure because i don't want to get it wrong

OpenStudy (anonymous):

oh well ill take that as a yes

OpenStudy (astrophysics):

It's right

OpenStudy (anonymous):

ok thanks @ganeshie8 @Astrophysics @Somy

OpenStudy (anonymous):

you were great help!!!

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

Now this will be a Chating room!!!

OpenStudy (anonymous):

chatting*

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

see ya suckas